A roller coaster (390 kg) moves from A (4.96 m above the ground) to B (20.9 m above the ground). Two non-conservative forces are present: friction does -1.86 x 104 J of work on the car, and a chain mechanism does +5.66 x 104 J of work to help the car up a long climb. What is the change in the car's kinetic energy KEf - KEo from A to B?
Answer #1

change in kinetic energy = -(change in potential energy) + work by non-conservative forces

change in potential energy = mass*g*change in height
change in potential energy = 390 * 9.81 * (20.9 - 4.96)
change in potential energy = 60984.846 J

work done by non-conservative forces = (5.66 - 1.86)*10^4
work done by non-conservative forces = 3.80*10^4 J

change in kinetic energy = - 60984.846 + 3.80*10^4


change in kinetic energy = - 22984.846 J